# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 3 Deep backward regression-based method
In this section, we formulate the DBR method to approximate the conditional expectations in scheme (2.4) using learned data. Our aim is to provide a fully implementable algorithm. Subsequently, we present a comprehensive error analysis of the DBR method. Finally, we investigate its convergence rate.
@@ 3.1 Deep learning method to approximate the conditional expectation
To implement $(Y_{i}^{\pi},Z_{i}^{\pi})$ based on the temporal discretization scheme (2.4), one has to approximate the conditional mathematics expectations. Now, we obtain them by means of the deep neural network method.
In what follows, the forward process $X_{\cdot}^{\pi,m}$ is simulated by Monte Carlo method in the following manners, for $i=0,1,\cdots,N-1$
$X_{i+1}^{\pi,m}=X_{i}^{\pi,m}+\mu(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m},\quad X_{0}^{\pi,m}=x_{0},$ (3.1)
@@
$\left\{\begin{aligned} Y_{i}^{\pi,m}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})\Big],\\ Z_{i}^{\pi,m}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\big].\end{aligned}\right$. (3.3)
@@
Now, the approximation algorithm is implemented in detail. The terms on the right hand side of (3.3) are written as
@@
by Monte Carlo method, where $Y_{i+1}^{\pi,m,k}=u(t_{i+1},X_{i+1}^{\pi,m,k})$ is independently and identically distributed as $Y_{i+1}^{\pi,m}$;
$R_{i,1}^{y}=$ $\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})\Big]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}-hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m}),$
$R_{i,1}^{z}=$ $\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right].$
Replacing the two random variables in local Monte Carlo approximates in (3.4) with two deep feedforward neural networks, namely
@@
$R_{i,2}^{y,0}=$ $\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{y,i}),$
$R_{i,2}^{z,0}=$ $\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right]-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{z,i}).$
From (3.4) and (3.5), we have
$\left\{\begin{aligned} Y_{i}^{\pi,m}=&\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{y,i})+R_{i,2}^{y,0}+R_{i,1}^{y},\\ Z_{i}^{\pi,m}=&\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{z,i})+R_{i,2}^{z,0}+R_{i,1}^{z}.\end{aligned}\right$. (3.6)
@@
$\left\{\begin{aligned} F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2},\\ F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2},\end{aligned}\right$. (3.7)
where $p_{i}^{m,k}=(t_{i},X_{i}^{\pi,m,k})$; $r_{j}\in\mathbb{N}^{+}$ denotes the number of iterations of the stochastic gradient descent (SGD) method; $\theta^{r_{j}}_{y,i}=\theta_{y,i}^{r_{j}-1}-\rho_{y,i}^{r_{j}-1}\nabla_{\theta_{y,i}^{r_{j}-1}}F_{y,i}^{r_{j}-1}(\theta_{y,i}^{r_{j}-1}),$ $\theta^{r_{j}}_{z,i}=\theta_{z,i}^{r_{j}-1}-\rho_{z,i}^{r_{j}-1}\nabla_{\theta_{z,i}^{r_{j}-1}}F_{z,i}^{r_{j}-1}(\theta_{z,i}^{r_{j}-1})$; $\rho_{y,i}^{r_{j}},\rho_{z,i}^{r_{j}}\in(0,+\infty)$ denote the learning rates; $\theta^{*}_{y,i}=\arg\min\limits_{\theta_{y,i}^{r_{j}}}F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})$, $\theta^{*}_{z,i}=\arg\min\limits_{\theta_{z,i}^{r_{j}}}F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})$;
$S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta_{y,i+1}^{*})\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right],$
@@
$\quad+hf(t_{i},X_{i}^{\pi,m},\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}}),\mathcal{NN}_{z,i+1,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})).$
Thus, by the optimization of the SGD method, the (3.6) is rewritten as
$\left\{\begin{aligned} Y_{i}^{\pi,m}=&\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i})+R_{i,2}^{y}+R_{i,1}^{y},\\ Z_{i}^{\pi,m}=&\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i})+R_{i,2}^{z}+R_{i,1}^{z},\end{aligned}\right$. (3.8)
@@
$R_{i,2}^{y}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*}),$
$R_{i,2}^{z}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right]-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}).$
Therefore, we deduce the fully discrete approximation solutions $(\mathcal{Y}_{i},\mathcal{Z}_{i})$ for the analytical solutions $(Y_{t_{i}},Z_{t_{i}})$ for $i=N,N-1,\cdots,1,0$, $m=1,2,\cdots,1,M$ based on the DBR method as below
@@
    the terminal condition is $(\mathcal{Y}_{N},\mathcal{Z}_{N})=(\Phi(X_{N}^{\pi,m}),\sigma^{\top}(t_{N},X^{\pi,m}_{N})D_{x}u(t_{N},X^{\pi,m}_{N}))$,
@@
    $\left\{\begin{aligned} \mathcal{Y}_{i}=&\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i}),\\ \mathcal{Z}_{i}=&\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i}).\end{aligned}\right$. (3.9)
@@
Input: Grid $0=t_{0}<\dots<t_{N}=T$ with step $h=T/N$.
Output: Trained networks $\{\mathcal{NN}_{y,i},\mathcal{NN}_{z,i}\}_{i=0}^{N-1}$ and learnable parameters.
Initialize: Networks $\theta_{y,i},\theta_{z,i}\sim$ Xavier initialization [18];
for $i=N-1,\dots,0$ do
 for $m=1,\dots,M$ do
      Generate $\Delta W_{i}^{m}\sim\mathcal{N}(0,h)$;
    $X^{\pi,m}_{i+1}\leftarrow X_{i}^{\pi,m}+b(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m}$, $X_{0}=x_{0}$;
   end for
  Generate $\Delta W_{i}^{m,k}\sim\mathcal{N}(0,h)$, for $m\in\{1,\cdots,M\},k\in\{1,\cdots,K\}$;
   Compute next states: $X_{i+1}^{\pi,m,k}\leftarrow X^{\pi,m}_{i}+b(t_{i},X^{\pi,m}_{i})h+\sigma(t_{i},X^{\pi,m}_{i})\Delta W_{i}^{m,k}$;
 if i==N1i==N-1 then
    $\mathcal{Y}_{i+1}^{m,k}\leftarrow g(t_{i+1},X_{i+1}^{\pi,m,k})$;
@@
    $\mathcal{Y}_{i+1}^{m,k}\leftarrow\mathcal{NN}_{y,i+1}(t_{i+1},X_{i+1}^{\pi,m,k},\theta^{*}_{y,i})$;
   end if
 $Loss_{Z,i}\leftarrow\frac{1}{M}\sum\limits_{m=1}^{M}\left[\left|\frac{1}{K}\sum\limits_{k=1}^{K}\left(\mathcal{Y}_{i+1}^{m,k}\frac{\Delta W_{i}^{m,k}}{h}\right)-\mathcal{NN}_{z,i}(t_{i},X_{i}^{\pi,m},\theta_{z,i})\right|^{2}\right]$;
 Update $\theta_{z,i}$ using the SGD method and the optimal parameter is denoted as $\theta^{*}_{z,i}$;
 $\mathcal{Z}_{i}^{m,k}\leftarrow\mathcal{NN}_{z,i}(t_{i},X_{i}^{\pi,m,k},\theta_{z,i}^{*})$;
 $\mathcal{Z}_{i}^{m}\leftarrow\mathcal{NN}_{z,i}(t_{i},X_{i}^{\pi,m},\theta_{z,i}^{*})$;
 $Loss_{Y,i}\leftarrow\frac{1}{M}\sum\limits_{m=1}^{M}\Bigg[\bigg|\frac{1}{K}\sum\limits_{k=1}^{K}\mathcal{Y}_{i+1}^{m,k}+f(t_{i},X_{i}^{\pi,m,k},\mathcal{NN}_{y,i}(t_{i},X_{i}^{\pi,m},\theta_{y,i}^{r}),\mathcal{Z}_{i}^{m,k})\cdot h$ $\mathcal{NN}_{y,i}(t_{i},X_{i}^{\pi,m},\theta_{y,i})\bigg|^{2}\Bigg]$;
 Update $\theta_{y,i}$ using the SGD method and the optimal parameter is denoted as $\theta^{*}_{y,i}$;
 $\mathcal{Y}_{i}^{m}\leftarrow\mathcal{NN}_{y,i}(t_{i},X_{i}^{\pi,m},\theta_{y,i}^{*})$;
end for
Final Estimation: $\mathcal{Y}_{0}=\mathcal{NN}_{y,0}(t_{0},X_{0},\theta^{*}_{y,0})$;

### 21 May 2026 v3

@@ 3 Deep backward regression-based method
In this section, we formulate the DBR method to approximate the conditional expectations in scheme (9) using learned data. Our aim is to provide a fully implementable algorithm. Subsequently, we present a comprehensive error analysis of the DBR method. Finally, we investigate its convergence rate.
@@ 3.1 Deep learning method to approximate the conditional expectation
To implement the time-discrete scheme (9), the conditional expectations must be approximated numerically. We approximate these conditional expectations by training neural networks at each time step.
In what follows, the forward process $X_{\cdot}^{\pi,m}$ is simulated by Monte Carlo simulation as follows, for $i=0,1,\cdots,N-1$
$X_{i+1}^{\pi,m}=X_{i}^{\pi,m}+\mu(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m},\quad X_{0}^{\pi,m}=x_{0},$ (11)
@@
$\left\{\begin{aligned} Y_{i}^{\pi,m}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})\Big],\\ Z_{i}^{\pi,m}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\big].\end{aligned}\right$. (13)
@@
Now, the approximation algorithm is implemented in detail. The terms on the right hand side of (13) are written as
@@
by Monte Carlo method, where $Y_{i+1}^{\pi,m,k}$ is independently and identically distributed as $Y_{i+1}^{\pi,m}$;
$R_{i,1}^{y}=$ $\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}\Big]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k},$
$R_{i,1}^{z}=$ $\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}.$
Consequently, following the Monte Carlo simulation, the computational formula for $(Y_{i}^{\pi,m},Z_{i}^{\pi,m})$ is given by
$\left\{\begin{aligned} Y_{i,K_{i}}^{\pi,m}&=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1,K_{i}}^{\pi,m,k}+hf\left(t_{i},X_{i}^{\pi,m},Y_{i,K_{i}}^{\pi,m},Z_{i,K_{i}}^{\pi,m}\right),\\ Z_{i,K_{i}}^{\pi,m}&=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1,K_{i}}^{\pi,m,k}\frac{(\Delta W_{i}^{m,k})^{\top}}{h}.\end{aligned}\right$. (15)
Replacing the two random variables in local Monte Carlo approximates in (15) with two deep feedforward neural networks, namely
@@
$R_{i,2}^{y,0}=$ $Y_{i,K_{i}}^{\pi,m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{y,i}),$
$R_{i,2}^{z,0}=$ $Z_{i,K_{i}}^{\pi,m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta^{0}_{z,i}).$
@@
$\left\{\begin{aligned} F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2},\\ F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2},\end{aligned}\right$. (17)
where $p_{i}^{m,k}=(t_{i},X_{i}^{\pi,m,k})$; $r_{j}\in\mathbb{N}^{+}$ denotes the number of iterations of the stochastic gradient descent (SGD) method; $\theta^{r_{j}}_{y,i}=\theta_{y,i}^{r_{j}-1}-\rho_{y,i}^{r_{j}-1}\nabla_{\theta_{y,i}^{r_{j}-1}}F_{y,i}^{r_{j}-1}(\theta_{y,i}^{r_{j}-1}),$ $\theta^{r_{j}}_{z,i}=\theta_{z,i}^{r_{j}-1}-\rho_{z,i}^{r_{j}-1}\nabla_{\theta_{z,i}^{r_{j}-1}}F_{z,i}^{r_{j}-1}(\theta_{z,i}^{r_{j}-1})$; $\rho_{y,i}^{r_{j}},\rho_{z,i}^{r_{j}}\in(0,+\infty)$ denote the learning rates; suppose that the SGD algorithm is capable of converging to the optimal solution: $\theta^{*}_{y,i}=\arg\min\limits_{\theta_{y,i}^{r_{j}}}F_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})$, $\theta^{*}_{z,i}=\arg\min\limits_{\theta_{z,i}^{r_{j}}}F_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})$;
$S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta_{y,i+1}^{*})\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h},$
@@
$\quad+hf(t_{i},X_{i}^{\pi,m},\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}}),\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})).$
Thus, by the optimization of the SGD method, (16) is rewritten as
$\left\{\begin{aligned} Y_{i,K_{i}}^{\pi,m}=&\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i})+R_{i,2}^{y},\\ Z_{i,K_{i}}^{\pi,m}=&\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i})+R_{i,2}^{z},\end{aligned}\right$. (18)
@@
$R_{i,2}^{y}$ $=Y_{i,K_{i}}^{\pi,m}-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*}),$
$R_{i,2}^{z}$ $=Z_{i,K_{i}}^{\pi,m}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}).$
Therefore, based on the DBR method, we define the fully discrete approximations $\{\mathcal{Y}_{i}\}_{i=0}^{N}$ and $\{\mathcal{Z}_{i}\}_{i=0}^{N-1}$ for the analytical solutions $(Y_{t_{i}},Z_{t_{i}})$, $i=0,1,\ldots,N$, $m=1,2,\ldots,M$, as follows:
@@
    the terminal condition is
    $\mathcal{Y}_{N}=g(X_{N}^{\pi,m}).$
    No terminal value $\mathcal{Z}_{N}$ is required by the backward scheme.
@@
    $\left\{\begin{aligned} \mathcal{Y}_{i}=&\ \mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i}),\\ \mathcal{Z}_{i}=&\ \mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i}).\end{aligned}\right$. (19)
For notational consistency in the subsequent error analysis, we use the following terminal-network convention:
$\mathcal{NN}_{y,N,1+d,n}^{\varrho,L,\mathbf{n}}(p_{N}^{m},\theta_{y,N}^{*})=g(X_{N}^{\pi,m}).$ (20)
Here $\mathcal{NN}_{y,N,1+d,n}^{\varrho,L,\mathbf{n}}$ is only a notational convention at the terminal time and no neural network is trained at $t_{N}$.
@@
Grid $0=t_{0}<t_{1}<\cdots<t_{N}=T$ with step $h=T/N$; the number $M$ of base Monte Carlo paths; the numbers $K_{i}$ of inner Monte Carlo samples. Trained networks $\{\mathcal{NN}_{y,i},\mathcal{NN}_{z,i}\}_{i=0}^{N-1}$ and the final estimate $\mathcal{Y}_{0}$.
Initialize the network parameters $\theta_{y,i},\theta_{z,i}$ for $i=0,\ldots,N-1$, e.g., by Xavier initialization;
Forward simulation of base paths:
for $m=1,\ldots,M$ do
   Set $X_{0}^{\pi,m}=x_{0}$;
 for $i=0,\ldots,N-1$ do
      Generate $\Delta W_{i}^{m}\sim N(0,hI_{d})$;
     Compute
$X_{i+1}^{\pi,m}\leftarrow X_{i}^{\pi,m}+\mu(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m}.$
Backward training:
for $i=N-1,\ldots,0$ do
 for $m=1,\ldots,M$ do
    for $k=1,\ldots,K_{i}$ do
         Generate $\Delta W_{i}^{m,k}\sim N(0,hI_{d})$;
        Compute the inner next state
$X_{i+1}^{\pi,m,k}\leftarrow X_{i}^{\pi,m}+\mu(t_{i},X_{i}^{\pi,m})h+\sigma(t_{i},X_{i}^{\pi,m})\Delta W_{i}^{m,k}.$
        Set
$p_{i}^{m}:=(t_{i},X_{i}^{\pi,m}),\qquad p_{i+1}^{m,k}:=(t_{i+1},X_{i+1}^{\pi,m,k}).$
       if i=N1i=N-1 then
$\mathcal{Y}_{i+1}^{m,k}\leftarrow g(X_{i+1}^{\pi,m,k});$
@@
$\mathcal{Y}_{i+1}^{m,k}\leftarrow\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m,k},\theta_{y,i+1}^{*});$
  Define the $Z$-loss
$\mathcal{L}_{z,i}(\theta_{z,i}):=\frac{1}{M}\sum_{m=1}^{M}\left|\frac{1}{K_{i}}\sum_{k=1}^{K_{i}}\mathcal{Y}_{i+1}^{m,k}\frac{(\Delta W_{i}^{m,k})^{\top}}{h}-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i})\right|^{2}.$
  Update $\theta_{z,i}$ by SGD, and denote the obtained optimal parameter by $\theta_{z,i}^{*}$;
  Define the $Y$-loss
$\mathcal{L}_{y,i}(\theta_{y,i}):=\frac{1}{M}\sum_{m=1}^{M}\Bigg|\frac{1}{K_{i}}\sum_{k=1}^{K_{i}}\mathcal{Y}_{i+1}^{m,k}+hf\left(t_{i},X_{i}^{\pi,m},\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}),\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\right)$
 $\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i})\Bigg|^{2}.$
  Update $\theta_{y,i}$ by SGD, and denote the obtained optimal parameter by $\theta_{y,i}^{*}$;
  Set, for $m=1,\ldots,M$,
$\mathcal{Y}_{i}^{m}\leftarrow\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*}),\qquad\mathcal{Z}_{i}^{m}\leftarrow\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}).$
Final estimation:
$\mathcal{Y}_{0}=\mathcal{NN}_{y,0,1+d,n}^{\varrho,L,\mathbf{n}}((t_{0},x_{0}),\theta_{y,0}^{*}).$